a bank account earns interest compounded continuously at a rate of 6% per year. how long would it take for the account to triple in value?
Answer by stanbon(57377)

a bank account earns interest compounded continuously at a rate of 6% per year. how long would it take for the account to triple in value?
---
A(t) = Ao*e^rt
---
Solve e^rt = 3
e^(0.06t) = 3
0.06t = ln(3)
t = 18.31 years
=====================
Cheers,
Stan H.
=====================